can someone help me pls
factor
x^3-4x^2-5x+20

Answer by Theo(13342)

Your equation is:

x^3-4x^2-5x+20

divide this equation by (x-4) to get:

(x-4) * (x^2-5)

This factors out to:

(x-4) * (x-sqrt(5)) * (x+sqrt(5))

I confirmed it's accurate by replacing x with 5 in both the original equation and the factored equation to get the same answer.

You can also multiply the factors out again to get back to the original equation.

I figured (x-4) might divide into the original equation because -4 * x^2 = -4x^2 which would subtract nicely from -4x^2 in the original equation, and becausse -4 * -5 = 20 which would subtract nicely from 20 at the end of the equation.

Not totally scientific but it worked.

x goes into x^3 by a factor of x^2

(x-4) * x^2 = x^3 - 4x^2 which subtracts from x^3 - 4x^2 without a remainder.

The division is left with (x-4) into (-5x + 20) which comes out very clean with a factor of -5.

-5 * (x-4) = -5x + 20 which subtracts very nicely from -5x + 20 to leave a remainder of 0.

My answer by dividing the original equation by (x-4) became (x-4) * (x^2 - 5) which led to the final solution.